Complete these ordered pairs for this equation.
(0, ), (2, ), (4, )3x + 2y = 6

Answer by Alan3354(69443)

Complete these ordered pairs for this equation.
(0, ), (2, ), (4, )
3x + 2y = 6
----------
All 3 have y missing.
3x + 2y = 6
Solve for y
y = -3x/2 + 3
---
(0, )
Sub 0 for x
y = 0 + 3 = 3
--> (0,3)
===========
(2, )
Sub 2 for x
y = -3*2/2 + 3 = 0
--> (2,0)
===========
(4, )
Same as above.